OpenStudy (anonymous):

Half angle identity for cos of 13pi /12 in radians or degree, for tan of 195 degrees , and cos of 19pi/12 in radians or degree whichever

OpenStudy (anonymous):

\(\frac{13\pi}{12}\) is half of \(\frac{13\pi}{6}\) use the half angle identity with \(\frac{13\pi}{6}\)

OpenStudy (anonymous):

\[\cos(\frac{13\pi}{12}=\pm\sqrt{\frac{1+\cos(\frac{13\pi}{6})}{2}}\]